OpenStudy (anonymous):

16 2/5% of 750 173 1/2-69 7/9 16 1/2 divided by 3/4 9 4/7 x 56

OpenStudy (anonymous):

1)16 2/5% = (5x16)+2/5% = 82/5x100 x 750 = 61500/500 = 123

OpenStudy (anonymous):

2) (173x2)+1/2 - (69x9)+7/9 = 347/2 - 628/9 = 3123/18 - 1256/18 = 1867/18 = 103 13/18

OpenStudy (anonymous):

3)(16x2)+1/2 x 4/3 = 33x4/3x2= 132/6= 22

OpenStudy (anonymous):

4) (9x7) + 4/7 x 56 = 67x56/7 = 536

OpenStudy (anonymous):

sorry i took long, hope it helps

OpenStudy (anonymous):

THANK YOU!!!